OpenStudy (anonymous):

Find the rate at which the volume of the sphere is changing when the radius of the sphere is 10cm and surface area is changing at Pi cm^2/min Find the rate at which the volume of the sphere is changing when the radius of the sphere is 10cm and surface area is changing at Pi cm^2/min @Mathematics

OpenStudy (anonymous):

This is what i did (is it right?) v=4/3 pi r^2 dv/dt = 4/3Pi*3r^2 dr/dt dv/dt = 4Pir^2*dr/dt

OpenStudy (anonymous):

i got 400Pi^2 cm^4/min

OpenStudy (anonymous):

Nevermind, I take that back. Here:

myininaya (myininaya):

\[V=\frac{4}{3} \pi r^3\] \[SA=4 \pi r^2\] --------------------- we want to find V' where r=10 and SA'=pi

OpenStudy (anonymous):

\[ V= \frac{4}{3} \pi r^3 \] \[\frac{dV}{dt} = 4\pi r^2 \frac{dr}{dt} \] but since \[ A = 4\pi r^2 \rightarrow \frac{dA}{dt} = 8 \pi r \frac{dr}{dt}\] which means \[\frac{dr}{dt} = \frac{1}{8\pi r}\frac{dA}{dt} \] and so \[\frac{dV}{dt} = 4 \pi r^2 \frac{1}{8 \pi r} \frac{dA}{dt} = \frac{r}{2} \frac{dA}{dt} = \frac{10cm}{2}\pi \space cm^2/min=5\pi\space cm^3/min\]

myininaya (myininaya):

\[V'=4 \pi r^2 r'=SA r'\] \[SA'=8 \pi r r'\] ------------------------------- \[V'_{r=10}=400 \pi r'=SA r'\] \[SA'_{r=10}=80 \pi r'=\pi => r'=\frac{1}{80} \] \[V'_{r=10}=SAr'=400 \pi \cdot \frac{1}{80} =5 \pi\] oh i'm slow

OpenStudy (anonymous):

thanks a lot friends! i wish i can buy you guys a beer =)